# A novel generalized inner product-based wave scattering from an underwater source in a compressible ocean

### v2

@@ 3 Solution technique
with $\mu_{0}(k)$ being the positive real root and $\mu_{n}(k)$ being of the form $\rm i\tilde{\mu}_{n}(k)$ for $n\geq 1$, where $\tilde{\mu}_{n}(k)$ are real, positive and in ascending order with respect to the index $n$. Note that while $\mu_{n}(k)$ depends on both $n$ and $k$, we will omit the argument $k$ when this is clear from context. The associated vertical eigenfunctions are written as
@@
where the prime denotes differentiation with respect to $t$. Using the following orthogonality results (which we prove in the following section):
@@
in which $D_{p}(k)$ are normalisation coefficients, $\delta(\cdot)$ is the Dirac delta and $\delta_{np}$ is the Kronecker delta, we obtain
@@
$\Phi(x,z,t)=\sum_{n=0}^{N}\int_{0}^{\infty}\left(\frac{\langle\Phi_{0}(x,z),\phi_{n}^{-}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{-}+\frac{\langle\Phi_{0}(x,z),\phi_{n}^{+}\rangle_{\mathcal{E}}}{D_{P}}\phi_{n}^{+}\right)\cos(\omega_{n}t)\,{\rm d}k\\ -\frac{1}{\rho}\sum_{n=0}^{N}\int_{0}^{\infty}\left(\frac{\langle P_{0}(x,z),\phi_{n}^{-}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{-}+\frac{\langle P_{0}(x,z),\phi_{n}^{+}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{+}\right)\frac{\sin(\omega_{n}t)}{\omega_{n}}\,{\rm d}k.$ (21)
4 Orthogonality of the eigenfunctions
Next, we determine the normalisation coefficients $D_{p}(k)$ while simultaneously confirming the orthogonality of the eigenfunctions. Firstly, we calculate the $L^{2}(-h,0)$ inner product of the depth-dependent functions, yielding
[…]
$\lim_{k\rightarrow\tilde{k}}\langle\phi_{n}^{\pm}(\cdot,\cdot,k),\phi_{p}^{\pm}(\cdot,\cdot,\tilde{k})\rangle_{\mathcal{E}}=\frac{c^{2}}{g}\left(\frac{\tilde{k}}{\mu_{n}\partial_{k}\tilde{\mu}_{n}}\right)\delta(k-\tilde{k})\delta_{np},$
where
For simplicity, we write
where
$D_{p}(k)=\left(\frac{h}{2}{\rm sech}^{2}{\mu_{p}h}+\frac{\tanh{\mu_{p}h}}{2\mu_{p}}+\frac{c^{2}}{g}\right).$
Similarly,
$\langle\phi_{n}^{\pm},\phi_{p}^{\mp}\rangle_{\mathcal{E}}=\frac{c^{2}}{g}\left(\frac{k^{2}-\tilde{k}^{2}}{\mu_{n}^{2}-\tilde{\mu}_{p}^{2}}\right)\delta(k+\tilde{k})=0,$ (23c)
since $k$ and $\tilde{k}$ are both positive real numbers.
@@ 5.1 Numerical evaluation
When evaluating (24), care must be taken due to the potentially highly oscillatory nature of the integrand (for large $x$ or $t$). We reformulate the integrals into the form
$I_{n}^{(1)}$ $=\sum_{\pm_{1}}\sum_{\pm_{2}}\pm_{2}\int_{0}^{\infty}g_{n\pm_{1}}^{(1)}e^{{\rm i}(\pm_{1}kx\pm_{2}\omega_{n}t)}\,{\rm d}k,$
$I_{n}^{(2)}$ $=\sum_{\pm_{1}}\sum_{\pm_{2}}\int_{0}^{\infty}g_{n\pm_{1}}^{(2)}e^{{\rm i}(\pm_{1}kx\pm_{2}\omega_{n}t)}\,{\rm d}k,$
@@
$g_{n\pm}^{(1)}=\frac{\omega_{n}\langle\Phi_{0}(x,z),\phi_{n}^{\pm}\rangle_{\mathcal{E}}f_{n}(z)}{2D_{p}}\quad\text{and}\quad g_{n\pm}^{(2)}=\frac{\langle P_{0}(x,z),\phi_{n}^{\pm}\rangle_{\mathcal{E}}f_{n}(z)}{2D_{p}}.$
Here, the subscripts of $\pm$ indicate whether the signs are chosen independently (different subscripts) or must be consistent (same subscript). This formulation allows us to implement a Filon-type quadrature method (Dahlquist & Björck, 2008; Filon, 1930) detailed in Appendix A. Further efficiency can be gained by recognising that
@@ 5.2 Underwater pressure disturbance
where $\hat{P}$ is the maximum pressure, $(x_{c},z_{c})$ is the centre of the pressure distribution and $\sigma$ is the size of the pressure. For the following results, we choose $\hat{P}=10^{6}$ Pa, $(x_{c},z_{c})=(0,-2000)$ m, and $\sigma=200$m.
@@ 6 Inclusion of static compression
In this section, we include (mild) density variance of the ocean water, the effect which is known in the literature as static compression, into the mathematical formulation, and develop the solution using a similar approach to that used in the previous section. The rigorous derivation of the governing equations can be found in the works of Longuet-Higgins (1950), Miyoshi (1954), and Sells (1965). Note that in the subsequent mathematical treatment, the density $\rho$ represents its value at the surface, which is treated as the baseline density. The hydrostatic density is $\rho e^{-\gamma z}$ (Abdolali & Kirby, 2017), where the effect of the variance in the density away from the surface is reflected through the term $\gamma$, which depends on the speed of sound in the water and is defined by $\gamma=g/c^{2}$. The effect of static compression is small for realistic oceans, as we shall see, and it significantly complicates the formulation. We believe that including it is worthwhile, first because it allows accurate calculation of its effect and second because it highlights how powerful the special inner product-based method is and how it can be extended to highly non-trivial cases. We note that the complexity arises in the theory, but that the computational difficulty is not greatly increased.
@@ 6.1 Mathematical Solution with Static Compression
$\iint_{\Omega}\left(\nabla^{2}\hat{\Phi}\,\hat{\Psi}-\hat{\Phi}\,\nabla^{2}\hat{\Psi}\right){\rm d}V=\oint_{\partial\Omega}\left(\hat{\Phi}_{n}\,\hat{\Psi}-\hat{\Phi}\,\hat{\Psi}_{n}\right){\rm d}S,$
Taking $\Omega$ to be the region $(-L,L)\times(-h,0)$ gives, in the limit as $L\to\infty$
@@
which is precisely in the form of (20). To finalise the mathematical solution in the case of static compression, it remains to compute the normalisation functions $D_{p}(k)$. This is done in Appendix B.
@@ 6.2 Time-domain simulations with static compression
$\frac{P_{\mathrm{static}}(x,z,t)-P(x,z,t)}{\max\limits_{x,z}(P_{\mathrm{static}}(x,z,t),P(x,z,t))},$
for $t=1$ (Figure 4), where $P_{\mathrm{static}}$ is the solution including static compression. We observe that the difference wave has a negative leading front in the upper half wave (i.e., upwards travelling waves) and a positive leading front in the lower half (i.e., downward travelling waves). We can see that the inclusion of static compression, on average, causes the magnitude of pressure above (below) the initial explosion location to decrease (increase) compared with the case with no static compression. It aligns with the fact that water is less (more) dense in the upper (lower) water region measured from the point of explosion when static compression is considered. There is a roughly $\pm 0.5\%$ pressure difference for the time considered ($0<t<20$) when including static compression. Finally, we note that the pressure wave speed has not changed with the addition of static compression. The Movie $3$ provides the full animation.
@@ 7 Conclusion
Animations for Figures 26 are uploaded as supplementary material (https://doi.org/10.6084/m9.figshare.30186271.v1).
@@
Appendix A Filon-type quadrature
@@
Appendix B Normalisation coefficients in the case of static compression
@@
$=\int_{-\infty}^{\infty}\langle f_{n}(z;k),f_{p}(z;\tilde{k})\rangle_{h}e^{\mp{\rm i}(k-\tilde{k})x}\,{\rm d}x=\frac{c^{2}}{g}\frac{\tilde{k}}{\tilde{\mu}_{n}\partial_{k}\tilde{\mu}_{n}}\delta(k-\tilde{k})\delta_{np}.$ (46a)
Note that in the case of counter-propagating waves, we compute $\langle\phi_{n}^{\pm}(\cdot,\cdot,k),\phi_{p}^{\mp}(\cdot,\cdot,\tilde{k})\rangle_{\mathcal{E}}=0$, which is analogous to (23c).

### v3

@@ 3 Solution technique
with $\mu_{0}(k)$ being the positive real root and $\mu_{n}(k)$ being of the form $\rm i\tilde{\mu}_{n}(k)$ for $n\geq 1$, where $\tilde{\mu}_{n}(k)$ are real, positive and in ascending order with respect to the index $n$. In addition, $k_{n}$ are real positive for all progressive acoustic-gravity wave modes and imaginary for evanescent modes. Note that while $\mu_{n}(k)$ depends on both $n$ and $k$, we will omit the argument $k$ when this is clear from context. The associated vertical eigenfunctions are written as
@@
where the prime denotes differentiation with respect to $t$. Using the following orthogonality results (which we prove in Appendix A):
@@
noting that (16a) is the inner product between waves travelling in the same direction and (16b) is the inner product between waves travelling in opposite directions. Here, in which $D_{p}(k)$ are normalisation coefficients, $\delta(\cdot)$ is the Dirac delta and $\delta_{np}$ is the Kronecker delta, we obtain
@@
$\Phi(x,z,t)=\sum_{n=0}^{\infty}\int_{0}^{\infty}\left(\frac{\langle\Phi_{0}(x,z),\phi_{n}^{-}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{-}+\frac{\langle\Phi_{0}(x,z),\phi_{n}^{+}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{+}\right)\cos(\omega_{n}t)\,{\rm d}k\\ -\frac{1}{\rho}\sum_{n=0}^{\infty}\int_{0}^{\infty}\left(\frac{\langle P_{0}(x,z),\phi_{n}^{-}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{-}+\frac{\langle P_{0}(x,z),\phi_{n}^{+}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{+}\right)\frac{\sin(\omega_{n}t)}{\omega_{n}}\,{\rm d}k.$ (23)
We note at this point some special features of this solution for $\Phi$. The solution can be calculated at any time from the initial conditions without reference to the solution at any other time. Moreover, the difficulty of solving for any time is equivalent as the method does not rely on temporal iteration. In particular, we compute the solution at $t=0$, which provides a validation of the method through comparison with the initial condition. We note that, provided that the $\phi_{n}^{\pm}$ are correct, then $\Phi$ must satisfy the governing equations, and if the potential matches at $t=0$, we have calculated the solution correctly.
@@ 4.1 Numerical evaluation
When evaluating (24), care must be taken due to the potentially highly oscillatory nature of the integrand (for large $x$ or $t$). We reformulate (23) into the form
$\Phi(x,z,t)=\sum_{n=0}^{\infty}I_{n}^{(1)}-\frac{1}{\rho}\sum_{n=0}^{\infty}I_{n}^{(2)},$
@@
$I_{n}^{(1)}$ $=\sum_{\pm}\sum_{j=\{0,1\}}(-1)^{j}\int_{0}^{\infty}g_{n\pm}^{(1)}e^{{\rm i}(\pm kx+(-1)^{j}\omega_{n}t)}\,{\rm d}k,$
$I_{n}^{(2)}$ $=\sum_{\pm}\sum_{j=\{0,1\}}\int_{0}^{\infty}g_{n\pm}^{(2)}e^{{\rm i}(\pm kx+(-1)^{j}\omega_{n}t)}\,{\rm d}k,$
and
$g_{n\pm}^{(1)}=\frac{\omega_{n}\langle\Phi_{0}(x,z),\phi_{n}^{\pm}\rangle_{\mathcal{E}}f_{n}(z)}{2D_{p}},\quad\quad g_{n\pm}^{(2)}=\frac{\langle P_{0}(x,z),\phi_{n}^{\pm}\rangle_{\mathcal{E}}f_{n}(z)}{2D_{p}}.$
The infinite sum is truncated at $N$ modes while the semi-infinite integral is truncated at $k=k_{\mathrm{max}}$ and discretised uniformly using spacing $\Delta k$. The choice of $N$, $k_{\mathrm{max}}$, and $\Delta k$ depends on the initial conditions and will be discussed in §4.2. In this paper, we take $N=100$ as that was sufficient to reproduce the initial condition. Similarly, the semi-infinite integral is truncated at $k=0.2$ and discretised uniformly using $1001$ points.
This formulation allows us to implement a Filon-type quadrature method (Dahlquist & Björck, 2008; Filon, 1930) detailed in Appendix B. Further efficiency can be gained by recognising that
@@ 4.2 Underwater pressure disturbance
where $\hat{P}$ is the maximum pressure, $(x_{c},z_{c})$ is the centre of the pressure distribution and $\sigma$ is the characteristic width of the pressure. For the following results, we choose $\hat{P}=10^{6}$ Pa, $(x_{c},z_{c})=(0,-2000)$ m, g=9.81g=9.81m/s2, c=1450c=1450m/s, h=4000h=4000m and $\sigma=200$m.
The numerical parameters $N$, $k_{\mathrm{max}}$, and $\Delta k$ are chosen by ensuring the initial condition can be reproduced with sufficient accuracy. We observe the maximum difference between the numerical solution and true initial condition over different parameter choices (see Appendix C for details). Based on this parameter sweep, we choose $N=100$, $k_{\mathrm{max}}=0.2$, and $\Delta k=0.0002$.
@@ 5 Inclusion of static compression
In this section, we include (mild) density variance of the ocean water, the effect which is known in the literature as static compression, into the mathematical formulation, and develop the solution using a similar approach to that used in the previous section. The rigorous derivation of the governing equations can be found in the works of Longuet-Higgins (1950), Miyoshi (1954), and Sells (1965). Note that in the subsequent mathematical treatment, the density $\rho$ represents its value at the surface, which is treated as the baseline density. The hydrostatic density is $\rho e^{-\gamma z}$ (Abdolali & Kirby, 2017), where the effect of the variance in the density away from the surface is reflected through the term $\gamma$, which depends on the speed of sound in the water and is defined by $\gamma=g/c^{2}$. This inclusion inherently modifies the underlying governing equation which we briefly show now.
Assuming $\rho_{0}(z)$ and $p_{0}(z)$ being the undisturbed water density and pressure, respectively, the fundamental equations of motion up to the first order yield (Bondi, 1947)
$\phi_{t}=-\int\frac{dp}{\rho}+gz,$ (27a)
$\rho_{t}+\nabla\cdot(\rho_{0}\nabla\phi)=0.$ (27b)
According to Longuet-Higgins (1950), pressure and density are related by the relationship
$\frac{\partial p}{\partial\rho}=c^{2},$ (28)
which modifies (27a) to
$\phi_{t}=-c^{2}\log\rho+gz.$ (29)
For the undisturbed state, one can obtain from (29)
$\rho_{0}=\hat{\rho}_{0}e^{gz/c^{2}}=\hat{\rho}_{0}e^{\gamma z},$ (30)
where $\hat{\rho}_{0}$ represents undisturbed density at $z=0$. Using $\rho=\rho_{0}(1+s)$, (29) and (27b) become
$\frac{\partial\phi}{\partial t}=-c^{2}s,$ (31a)
$\frac{\partial s}{\partial t}+\nabla^{2}\phi+\gamma\phi_{z}=0.$ (31b)
Eliminating $s$ yields the governing equation
$\nabla^{2}\Phi=\frac{1}{c^{2}}\Phi_{tt}+\gamma\Phi_{z}.$ (32)
The effect of static compression is small for realistic oceans (Abdolali et al., 2019), as we shall see, and it significantly complicates the formulation. We believe that including it is worthwhile, first because it allows accurate calculation of its effect and second because it highlights how powerful the special inner product-based method is and how it can be extended to highly non-trivial cases. We note that the complexity arises in the theory, but that the computational difficulty is not greatly increased.
@@ 5.1 Mathematical Solution with Static Compression
$\iint_{\Omega}\left(\nabla^{2}\hat{\Phi}\,\hat{\Psi}-\hat{\Phi}\,\nabla^{2}\hat{\Psi}\right){\rm d}V=\oint_{\partial\Omega}\left(\hat{\Phi}_{n}\,\hat{\Psi}-\hat{\Phi}\,\hat{\Psi}_{n}\right){\rm d}S.$
The choices of $\hat{\Phi}$ and $\hat{\Psi}$ ensure elimination of the first order derivative term in (33a). Now taking $\Omega$ to be the region $(-L,L)\times(-h,0)$ gives, in the limit as $L\to\infty$
@@
which is precisely in the form of (22). To finalise the mathematical solution in the case of static compression, it remains to compute the normalisation functions $D_{p}(k)$. This is done in Appendix D. We note that (46) and (47) reduce to (23) and (22), respectively, by setting $\gamma=0$.
@@ 5.2 Time-domain simulations with static compression
$\frac{P_{\mathrm{static}}(x,z,t)-P(x,z,t)}{\max\limits_{x,z}(P_{\mathrm{static}}(x,z,t),P(x,z,t))},$ (48)
for $t=1$ (Figure 4), where $P_{\mathrm{static}}$ is the solution including static compression. We observe that the difference wave has a negative leading front in the upper half wave (i.e., upwards travelling waves) and a positive leading front in the lower half (i.e., downward travelling waves). We can see that the inclusion of static compression, on average, causes the magnitude of pressure above (below) the initial explosion location to decrease (increase) compared with the case with no static compression. It aligns with the fact that water is less (more) dense in the upper (lower) water region measured from the point of explosion when static compression is considered.
@@
Further, we calculated the maximum of the percentage difference (48) over space ($|x|<10^{5}$,h<z<0-h<z<0) and time ($0<t<20$) for different depths, $h$, where the pressure is centred halfway between the bottom and the surface $(x_{c},z_{c})=(0,-h/2)$ and for a pressure centre fixed at (0,250)(0,-250). From the results presented in Table 1, we can see that the pressure difference increases as the depth decreases, and in the case of the pressure at $z_{c}=-250$m, there is a minimum in pressure difference due to static compression at roughly h=1500h=1500m before the difference increases as the depth increases. However, the differences still remain roughly within $\pm 1\%$ for all tested depths. Finally, we note that the pressure wave speed has not changed with the addition of static compression. The Movie $3$ provides the full animation for Figure 4.
Depth ($h$) Percentage pressure Percentage pressure
difference for difference for
$z_{c}=-h/2$ (%) $z_{c}=-250$ (%)
500 1.2415 1.2415
1000 0.6565 0.6688
1500 0.5042 0.5877
2000 0.4572 0.6535
2500 0.5306 0.6978
3000 0.4935 0.8619
3500 0.5170 0.8744
4000 0.5988 1.0047
Table 1: The maximum percentage pressure difference (48) for different depths, $h$, where the pressure is centred either halfway between the bottom and the surface $(x_{c},z_{c})=(0,-h/2)$, or at (0,250)(0,-250).
@@ 6 Conclusion
Animations for Figures 26 are uploaded as supplementary material.
@@
Appendix A Orthogonality of the eigenfunctions
Here, we determine the normalisation coefficients $D_{p}(k)$ while simultaneously confirming the orthogonality of the eigenfunctions. Firstly, we calculate the $L^{2}(-h,0)$ inner product of the depth-dependent functions through two applications of integration by parts, yielding
[…]
$=\int_{-\infty}^{\infty}\left(\frac{\omega_{n}^{2}-\tilde{\omega}_{p}^{2}}{\mu_{n}^{2}-\tilde{\mu}_{p}^{2}}+\frac{c^{2}}{g}\right)e^{\mp{\rm i}(k-\tilde{k})x}\,{\rm d}x,$
$=\frac{c^{2}}{g}\int_{-\infty}^{\infty}\left(\frac{k^{2}-\tilde{k}^{2}}{\mu_{n}^{2}-\tilde{\mu}_{p}^{2}}\right)e^{\mp{\rm i}(k-\tilde{k})x}\,{\rm d}x,$
[…]
$\lim_{k\rightarrow\tilde{k}}\langle\phi_{n}^{\pm}(\cdot,\cdot,k),\phi_{p}^{\pm}(\cdot,\cdot,\tilde{k})\rangle_{\mathcal{E}}=2\pi\frac{c^{2}}{g}\left(\frac{\tilde{k}}{\mu_{n}\partial_{k}\tilde{\mu}_{n}}\right)\delta(k-\tilde{k})\delta_{np},$
where
For simplicity, we write
where
$D_{p}(k)=2\pi\left(\frac{h}{2}{\rm sech}^{2}{\mu_{p}h}+\frac{\tanh{\mu_{p}h}}{2\mu_{p}}+\frac{c^{2}}{g}\right).$
Similarly,
$\langle\phi_{n}^{\pm},\phi_{p}^{\mp}\rangle_{\mathcal{E}}=2\pi\frac{c^{2}}{g}\left(\frac{k^{2}-\tilde{k}^{2}}{\mu_{n}^{2}-\tilde{\mu}_{p}^{2}}\right)\delta(k+\tilde{k})=0,$ (51b)
since $k$ and $\tilde{k}$ are both positive real numbers.
Appendix B Filon-type quadrature
@@
Appendix C Errors for numerical parameter choice
This section compares the effect that the numerical scheme parameters $N$, $k_{\mathrm{max}}$, and $\Delta k$ have on the error in reconstructing the initial condition for the underwater Gaussian pressure disturbance (26). The physical parameters used are $\hat{P}=10^{6}$ Pa, $(x_{c},z_{c})=(0,-2000)$ m, g=9.81g=9.81m/s2, c=1450c=1450m/s, h=4000h=4000m and $\sigma=200$m. The error measure is given by
$\max_{x,z}\frac{\rho\Phi_{t}(x,z,0)+P_{0}(x,y)}{\hat{P}}$,
as the strength of the pressure $\hat{P}$ acts as a linear scaling of the solution.
Tables 2-4 present the errors when varying $N$, $k_{\mathrm{max}}$, and $\Delta k$, respectively. When not varied, the default parameter values $N=100$, $k_{\mathrm{max}}=0.2$, and $\Delta k=0.0002$ are used. We note that the error given by the default parameters is sufficient for visual convergence. Of the three parameters, the number of included modes $N$ (Table 2) has the most significant effect on the accuracy of the solution because it determines how well the initial condition can be resolved in the vertical direction. The choice of $k_{\mathrm{max}}$ (Table 3) is affected by the choice of pressure distribution, in our case, leading to exponential decay in the integrand. Finally, all tested discretisation spacings $\Delta k$ (Table 4) resulted in similar error, indicating that we have sufficient resolution of our integrand to numerically perform the integration.
NN Error
[…]
Table 2: Initial condition reconstruction error for different NN, and $k_{\mathrm{max}}=0.2$ and $\Delta k=0.0002$.
$k_{\mathrm{max}}$ Error
[…]
Table 3: Initial condition reconstruction error for different $k_{\mathrm{max}}$, and $N=100$ and $\Delta k=0.0002$.
$\Delta k$ Error
[…]
Table 4: Initial condition reconstruction error for different $\Delta k$, and $k_{\mathrm{max}}=0.2$ and $N=100$.
Appendix D Normalisation coefficients in the case of static compression
@@
$=\int_{-\infty}^{\infty}\langle f_{n}(z;k),f_{p}(z;\tilde{k})\rangle_{h}e^{\mp{\rm i}(k-\tilde{k})x}\,{\rm d}x=2\pi\frac{c^{2}}{g}\frac{\tilde{k}}{\tilde{\mu}_{n}\partial_{k}\tilde{\mu}_{n}}\delta(k-\tilde{k})\delta_{np}.$ (55a)
Note that in the case of counter-propagating waves, we compute $\langle\phi_{n}^{\pm}(\cdot,\cdot,k),\phi_{p}^{\mp}(\cdot,\cdot,\tilde{k})\rangle_{\mathcal{E}}=0$, which is analogous to (51b).